# On the Stability of Graph Convolutional Neural Networks: A Probabilistic Perspective

### v3

@@ 1 Introduction
A probabilistic formulation (Section 3). We propose a novel probabilistic framework for analyzing the embedding stability of graph filters and GCNNs under edge perturbation. Unlike prior work that focuses on the worst-case scenarios, we introduce the notion of expected embedding perturbation, enabling a more representative assessment of stability across a broad range of input graph signals. Under this framework, we derive an exact characterization of the stability of graph filters and an upper bound for multilayer GCNNs. Both results are distribution-aware, revealing how embedding deviations are jointly determined by the second-moment matrix of graph signals and graph topology perturbations. To the best of our knowledge, this is the first work to establish a principled framework for understanding the interplay between data distribution and structural perturbations in the context of embedding stability.
@@
Related work. Analyzing the inference-time stability of GCNNs has gained growing interest in recent years due to its popularity in a variety of real-world applications. Existing theoretical studies investigate different types of perturbations: Gao et al. [18], Testa et al. [40, 41], Wang et al. [46], Ceci and Barbarossa [7] study the embedding perturbation induced by stochastically adding or deleting a small number of edges. In Liao et al. [31], the authors analyzed the change in embeddings of GCNNs under a perturbation applied to the learnable model parameters. Ruiz et al. [36] adopted a relative perturbation model where the perturbation on the graph is formulated as a multiplicative factor applied to the graph shift operator. Keriven et al. [25] studies the stability of GCNNs to small deformations of the random graph model. Our work is most closely related to a series of studies on the embedding stability under deterministic edge perturbation on the graph [28, 29, 22, 24, 16, 34]. The main difference between our work and existing embedding stability studies is that prior analyses are limited to the worst-case embedding perturbations, while our work first explores a more representative notion using expected embedding perturbations. Such formulation not only captures embedding behaviour from a broader range of inputs, but also enables a principled integration of input distribution into the stability analysis.
@@ 2.1 Notation and definitions
Graph Filters. The structural information of a graph $\mathcal{G}$ is encoded via a graph shift operator $\mathbf{S}$, a self-adjoint matrix satisfying $\mathbf{S}_{i,j}=0$ for all $(i,j)\notin\mathcal{E}$. Common choices for $\mathcal{S}$ include the graph adjacency matrix $\mathbf{A}$ and the graph Laplacian $\mathbf{L}$ [5]. A graph filter is defined as a function of the graph shift operator, denoted as $g(\mathbf{S})$. Typical examples of graph filters include polynomial filters, $g(\mathbf{S})=\sum_{k=0}^{K}c_{k}\mathbf{S}^{k}$ and autoregressive moving average filter [20]. Given input graph signals $\mathbf{X}$, the filter output (i.e., output embedding) is expressed as $g(\mathbf{S})\mathbf{X}$.
@@ 2.2 Problem formulation
To maintain focus, we do not explicitly address the permutation equivariance of graph filters and GCNNs, which is considered in other stability studies such as Gama et al. [16]. However, our framework can readily incorporate such equivariance by only replacing the filter perturbation $\mathbf{E}_{g}=g(\mathbf{S})-g(\mathbf{S}_{p})$ with its permutation-modulated counterpart $\mathbf{E}_{g}=g(\mathbf{S})-g(\mathbf{P}_{0}\mathbf{S}_{p}\mathbf{P}_{0}^{T})$ where $\mathbf{P}_{0}=\operatorname*{arg\,min}_{\mathbf{P}\in\Pi_{n}}{\|\mathbf{P}^{T}\mathbf{S}\mathbf{P}-\mathbf{S}_{p}\|}$.
@@ 3.1 Stability Analysis of Graph Filters and Single-Layer GCNNs
Let $\sigma(\cdot)$ be a monotonically non-decreasing activation function. Then the expected embedding perturbation of a single-layer GCNN is monotonically non-decreasing in the expected embedding perturbation of its associated graph filter:
@@ 3.2 Stability Analysis of Multilayer GCNNs
$\mathbb{E}_{X\sim\mathcal{D}}\left[\|{\mathbf{X}}^{(L)}-{\mathbf{X}}_{p}^{(L)}\|_{F}^{2}\right]\leq{d}\underbrace{C_{\sigma}^{2L}C^{2L-2}\prod_{j=1}^{L}\|\mathbf{\Theta}^{(j)}\|^{2}}_{{\text{model}}}\Big{(}\underbrace{(L-1)}_{\textit{model}}\underbrace{\|\mathbf{E}_{g}\|^{2}}_{\textit{pert.}}\underbrace{\mathrm{tr}(\mathbf{K})}_{\textit{signals}}+\underbrace{\langle\mathbf{K},\mathbf{E}_{g}^{T}\mathbf{E}_{g}\rangle}_{\textit{K-modulated pert.}}\Big{)}.$ (5)
Our result in (5) shows that the expected embedding perturbation of an $L$-layer GCNN is governed by several key components. The first is model complexity, captured by the factor $C_{\sigma}^{2L}C^{2L-2}\prod_{j=1}^{L}\|\mathbf{\Theta}^{(j)}\|^{2}$ and $(L-1)$, which reflects the cumulative effects of activation nonlinearity, filter norms, and learned weights. The bound also depends on the filter perturbation norm $\|\mathbf{E}_{g}\|^{2}$ and trace of the second-moment matrix $\mathrm{Tr}(\mathbf{K})$, which reflect the magnitudes of the filter change and input intensity, respectively. The second-moment matrix modulated filter perturbation term $\langle\mathbf{K},\mathbf{E}_{g}^{T}\mathbf{E}_{g}\rangle$, is derived from bounding the recursive propagation of perturbations across layers. This term highlights how the embedding change is shaped jointly by the graph filter perturbation and the second-order statistics of the graph signal. For further intuition and a complete derivation, we refer the reader to the proof in Appendix B.4.
In prior work, Kenlay et al. [23] showed that for unit-length input signals, the worst-case embedding perturbation of an $L$-layer GCN [26] satisfies $\sup\|\mathbf{x}^{(L)}-\mathbf{x}_{p}^{(L)}\|_{F}^{2}\leq L^{2}\prod_{j=1}^{L}\|\mathbf{\Theta}^{(j)}\|^{2}d\|\mathbf{E}_{g}\|^{2}$. Compared to this worst-case bound, our result presents a tighter characterization of the embedding stability. More importantly, our result is distribution-aware, capturing the joint influence of second-order statistics and graph topology perturbations on the embedding deviation.
@@ 4 A Structural Interpretation
$\mathbb{E}_{X\sim\mathcal{D}}[\|\mathbf{E}_{g}X\|_{2}^{2}]=\langle\mathbf{K},\mathbf{E}_{g}^{T}\mathbf{E}_{g}\rangle=\sum_{\{u,v\}\in\mathcal{P}}(\mathbf{K}_{uu}+\mathbf{K}_{vv})+2\sum_{\begin{subarray}{c}\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}\end{subarray}}\sigma_{uv}\sigma_{uv^{\prime}}\mathbf{K}_{vv^{\prime}}.$ (6)
@@
$\mathbb{E}_{X\sim\mathcal{D}}[\|\mathbf{E}_{g}X\|_{2}^{2}]=\langle\mathbf{K},\mathbf{E}_{g}^{T}\mathbf{E}_{g}\rangle=2\sum_{\{u,v\}\in\mathcal{P}}\mathcal{R}(u,v)+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{uv^{\prime}}(\mathcal{R}(u,v)+\mathcal{R}(u,v^{\prime})-\mathcal{R}(v,v^{\prime})),$ (7)
@@
Several prior studies [30, 43, 55] have shown that certain structural patterns of adversarial edge perturbations are more impactful than others. The key distinction between these studies and our analysis lies in the scope of consideration: previous approaches design perturbations primarily with respect to the graph topology, such as vertex degree, whereas our framework jointly accounts for both the graph topology and the signal distribution. In earlier work, Waniek et al. [47] proposed the heuristic perturbation method DICE, which aims to disconnect nodes within the same community and connect nodes across communities. Our theoretical analysis, along with the cSBM case study, provides a rigorous justification for this heuristic by explaining from a perspective of embedding perturbation why and when this strategy is efficient. We emphasize that our framework is broadly applicable, as it does not rely on a specific graph filter nor require ground-truth labels from downstream tasks. This generality moves beyond prior heuristics and provides principled insights into the structure of impactful perturbations across diverse graph learning settings.
@@ 5 Empirical Study on Adversarial Edge Perturbation
$\mathcal{P}^{*}=\operatorname*{arg\,max}_{|\mathcal{P}|=m}\mathbb{E}[\|\mathbf{E}_{g}X\|_{F}^{2}],\;\ \text{or }\mathcal{P}^{*}=\operatorname*{arg\,max}_{|\mathcal{P}|=m}\mathbb{E}[\|X^{(L)}-X^{(L)}_{p}\|_{F}^{2}].\vspace{0pt}$
@@
$\hat{\mathcal{P}}=\operatorname*{arg\,max}_{|\mathcal{P}|=m}\langle\mathbf{K},\mathbf{E}_{g}^{T}\mathbf{E}_{g}\rangle.\vspace{0pt}$
Here, the optimization objective involves the filter perturbation $\mathbf{E}_{g}$ and second-moment matrix $\mathbf{K}$. The filter perturbation $\mathbf{E}_{g}$ can be computed directly from its definition, $\mathbf{E}_{g}=g(\mathbf{S})-g(\mathbf{S}_{p})$, for any given filter and perturbation $\mathcal{P}$ (examples are provided in Appendix A.2). The second-moment matrix can be obtained using the empirical second-moment, i.e., $\mathbf{K}_{\text{samp}}=1/N\sum_{i=1}^{N}\mathbf{x}_{i}\mathbf{x}_{i}^{T}$, with alternative approaches discussed in Appendix C.3. The main obstacle to solving this optimization problem is the computational cost, as the complexity of brute-force searching is $O\left(n^{m}\right)$. To circumvent this, we consider relaxing discrete constraints in the problem and applying a projected gradient descent method developed by Xu et al. [51] for efficient adversarial attack in Prob-PGD. We present the algorithmic details in Appendix C.1.
@@ 5.1 Adversarial embedding attack
Graph signals. Among the above datasets, only the ENZYMES dataset contains real-world measurement signals associated with their vertices, while the remaining five datasets only contain graph structures without signals on their vertices. For these graph-only datasets, we generate 100 synthetic signals from an arbitrary distribution, such as random Gaussian signals, signals from cSBM, or the smoothness distribution (see Appendix C.2 for a detailed description). The second-moment matrix $\mathbf{K}$ is obtained from computing the sample graph signals, i.e, using $\mathbf{K}_{\text{samp}}=\sum_{i=1}^{100}\mathbf{x}_{i}\mathbf{x}_{i}^{T}$.
@@ 5.2 Adversarial attack on downstream tasks
Graph classification. We evaluate classification performance on two benchmark datasets: MUTAG (188 molecular graphs divided into two classes) and ENZYMES (600 protein graphs labeled by six enzyme classes). Both datasets provide numerical node features (e.g., atom types and physical measurements), from which we compute the second-moment matrix as $\mathbf{K}_{\text{samp}}=\sum_{i=1}^{d}\mathbf{x}_{i}\mathbf{x}_{i}^{T}$. For each dataset, we use 80% of the graphs to pre-train 10 GCNNs and keep their parameters fixed during evaluation. To assess their prediction robustness, we perturb $5\%$ edges in test set graphs using different attack algorithms and measure classification accuracy before and after perturbation.
Node classification. We use the Cora citation network (2708 vertices belonging to seven categories) with sparse bag-of-words node features. We pre-train 10 GCNNs on the unperturbed graphs with 140 labeled nodes and keep the model parameters fixed during evaluation. To assess the robustness of these pre-trained models, we apply different algorithms to perturb 1000 edges to the graph and evaluate classification accuracy on test set nodes before and after perturbation.
@@
MUTAG Unpert. 70.00±$\pm 2.68$ 82.89±$\pm 4.60$ 68.95±$\pm 1.58$
Random 73.16±$\pm 1.05$ 70.53±$\pm 4.82$ 72.89±$\pm 1.21$ 2.24±$\pm 0.15$ 28.62±$\pm 2.34$ 3.57±$\pm 0.17$
Wst-PGD 69.21±$\pm 2.37$ 76.32±$\pm 8.15$ 68.16±$\pm 1.42$ 1.95±$\pm 0.13$ 37.47±$\pm 9.47$ 3.11±$\pm 0.16$
Prob-PGD 42.89±$\pm 4.86$ 76.58±$\pm 8.02$ 41.58±$\pm 3.87$ 3.09±$\pm 0.16$ 96.26±$\pm 11.39$ 5.21±$\pm 0.25$
ENZYMES Unpert. 56.33±$\pm 3.10$ 56.75±$\pm 2.99$ 54.75±$\pm 3.48$
Random 49.58±$\pm 3.93$ 45.83±$\pm 2.64$ 45.50±$\pm 2.94$ 15.3±$\pm 0.7$ 81.3±$\pm 11.3$ 36.4±$\pm 1.2$
Wst-PGD 52.58±$\pm 3.56$ 39.75±$\pm 3.29$ 48.75±$\pm 4.27$ 11.5±$\pm 0.5$ 264.9±$\pm 34.7$ 27.2±$\pm 0.9$
Prob-PGD 44.25±$\pm 2.09$ 36.75±$\pm 4.62$ 40.75±$\pm 3.83$ 17.9±$\pm 0.9$ 297.9±$\pm 57.1$ 43.6±$\pm 1.6$
Cora Unpert. 80.06±$\pm 0.51$ 74.92±$\pm 1.86$ 80.33±$\pm 0.29$
Random 78.30±$\pm 0.82$ 71.75±$\pm 1.85$ 78.12±$\pm 0.64$ 93±$\pm 4$ 974±$\pm 496$ 136±$\pm 4$
Wst-PGD 78.47±$\pm 0.44$ 73.76±$\pm 2.07$ 78.30±$\pm 0.33$ 97±$\pm 7$ 16901±$\pm 12495$ 151±$\pm 5$
Prob-PGD 77.88±$\pm 0.33$ 56.67±$\pm 4.57$ 77.49±$\pm 0.29$ 106±$\pm 5$ 43565±$\pm 33035$ 160±$\pm 6$

### v4

@@ 1 Introduction
A probabilistic formulation (Section 3). We propose a probabilistic framework for analyzing the embedding stability of graph filters and GCNNs under edge perturbation. Unlike prior work that focuses on the worst-case scenarios, we introduce the notion of expected embedding perturbation, enabling a more representative assessment of stability across a broad range of input graph signals. Under this framework, we derive an exact characterization of the stability of graph filters and an upper bound for multilayer GCNNs. Both results are distribution-aware, revealing how embedding deviations are jointly determined by the second-moment matrix of graph signals and graph topology perturbations. To the best of our knowledge, this is the first work to establish a framework for understanding the interplay between data distribution and structural perturbations in the context of embedding stability.
@@
Related work. Analyzing the inference-time stability of GCNNs has gained growing interest in recent years due to its popularity in a variety of real-world applications. Existing theoretical studies investigate different types of perturbations: Gao et al. [18], Testa et al. [43, 44], Wang et al. [49], Ceci and Barbarossa [7] study the embedding perturbation induced by stochastically adding or deleting a small number of edges. In Liao et al. [32], the authors analyzed the change in embeddings of GCNNs under a perturbation applied to the learnable model parameters. Ruiz et al. [39] adopted a relative perturbation model where the perturbation on the graph is formulated as a multiplicative factor applied to the graph shift operator. Keriven et al. [26] studies the stability of GCNNs to small deformations of the random graph model. Our work is most closely related to a series of studies on the embedding stability under deterministic edge perturbation on the graph [29, 30, 23, 25, 16, 35, 37, 38]. The main difference between our work and existing embedding stability studies is that prior analyses are limited to the worst-case embedding perturbations, while our work first explores a more representative notion using expected embedding perturbations. Such formulation not only captures embedding behaviour from a broader range of inputs, but also enables the integration of input distribution into the stability analysis.
@@ 2.1 Notation and definitions
Graph Filters. The structural information of a graph $\mathcal{G}$ is encoded via a graph shift operator $\mathbf{S}$, a self-adjoint matrix satisfying $\mathbf{S}_{i,j}=0$ for all $(i,j)\notin\mathcal{E}$. Common choices for $\mathbf{S}$ include the graph adjacency matrix $\mathbf{A}$ and the graph Laplacian $\mathbf{L}$ [5]. A graph filter is defined as a function of the graph shift operator, denoted as $g(\mathbf{S})$. Typical examples of graph filters include polynomial filters, $g(\mathbf{S})=\sum_{k=0}^{K}c_{k}\mathbf{S}^{k}$ and autoregressive moving average filter [21]. Given input graph signals $\mathbf{X}$, the filter output (i.e., output embedding) is expressed as $g(\mathbf{S})\mathbf{X}$.
@@ 2.2 Problem formulation
To maintain focus, we do not explicitly address the permutation equivariance of graph filters and GCNNs, which is considered in other stability studies such as Gama et al. [16]. However, our framework can readily incorporate such equivariance by only replacing the filter perturbation $\mathbf{E}_{g}=g(\mathbf{S})-g(\mathbf{S}_{p})$ with its permutation-modulated counterpart $\mathbf{E}_{g}=g(\mathbf{S})-g(\mathbf{P}_{0}\mathbf{S}_{p}\mathbf{P}_{0}^{T})$ where $\mathbf{P}_{0}=\operatorname*{arg\,min}_{\mathbf{P}\in\Pi_{n}}{\|\mathbf{P}^{T}\mathbf{S}\mathbf{P}-\mathbf{S}_{p}\|_{F}}$.
@@ 3.1 Stability Analysis of Graph Filters and Single-Layer GCNNs
Let $\sigma(\cdot)$ be a monotonically non-decreasing activation function. Then the expected embedding perturbation of a single-layer GCNN is monotonically non-decreasing in the expected embedding perturbation of its associated graph filter.
@@ 3.2 Stability Analysis of Multilayer GCNNs
$\mathbb{E}_{X\sim\mathcal{D}}\hskip-2.0pt\left[\|{\mathbf{X}}^{(L)}-{\mathbf{X}}_{p}^{(L)}\|_{F}^{2}\right]\hskip-1.0pt\leq\hskip-1.0pt{d}\underbrace{C_{\sigma}^{2L}C^{2L-2}\prod_{j=1}^{L}\|\mathbf{\Theta}^{(j)}\|^{2}}_{{\text{model}}}\Big(\hskip-4.0pt\underbrace{(L-1)}_{\textit{model}}\underbrace{\|\mathbf{E}_{g}\|^{2}}_{\textit{pert.}}\underbrace{\mathrm{tr}(\mathbf{K})}_{\textit{signals}}\hskip-2.0pt+\hskip-3.0pt\underbrace{\langle\mathbf{K},\mathbf{E}_{g}^{T}\mathbf{E}_{g}\rangle}_{\textit{K-modulated pert.}}\hskip-6.0pt\Big).$ (5)
Our result in (5) shows that the expected embedding perturbation of an $L$-layer GCNN is governed by several key components. The first is model complexity, captured by the factor $C_{\sigma}^{2L}C^{2L-2}\prod_{j=1}^{L}\|\mathbf{\Theta}^{(j)}\|^{2}$ and (L1)(L-1), which reflects the cumulative effects of activation nonlinearity, filter norms, and learned weights. The bound also depends on the filter perturbation norm $\|\mathbf{E}_{g}\|^{2}$ and trace of the second-moment matrix $\mathrm{Tr}(\mathbf{K})$, which reflect the magnitudes of the filter change and input intensity, respectively. The second-moment matrix modulated filter perturbation term $\langle\mathbf{K},\mathbf{E}_{g}^{T}\mathbf{E}_{g}\rangle$, is derived from bounding the recursive propagation of perturbations across layers. This term highlights how the embedding change is shaped jointly by the graph filter perturbation and the second-order statistics of the graph signal. For further intuition and a complete derivation, we refer the reader to the proof in Appendix B.4.
In prior work, Kenlay et al. [24] showed that for unit-length input signals, the worst-case embedding perturbation of an $L$-layer GCN [27] satisfies $\sup\|\mathbf{x}^{(L)}-\mathbf{x}_{p}^{(L)}\|_{F}^{2}\leq dL^{2}\prod_{j=1}^{L}\|\mathbf{\Theta}^{(j)}\|^{2}\|\mathbf{E}_{g}\|^{2}$. Compared to this worst-case bound, our result presents a tighter characterization of the embedding stability. More importantly, our result is distribution-aware, capturing the joint influence of second-order statistics and graph topology perturbations on the embedding deviation.
@@ 4 A Structural Interpretation
$\mathbb{E}_{X\sim\mathcal{D}}[\|\mathbf{E}_{g}X\|_{2}^{2}]=\langle\mathbf{K},\mathbf{E}_{g}^{T}\mathbf{E}_{g}\rangle=\sum_{\{u,v\}\in\mathcal{P}}(\mathbf{K}_{uu}+\mathbf{K}_{vv})+2\hskip-12.0pt\sum_{\begin{subarray}{c}\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}\end{subarray}}\sigma_{uv}\sigma_{uv^{\prime}}\mathbf{K}_{vv^{\prime}}.$ (6)
@@
$\mathbb{E}_{X\sim\mathcal{D}}[\|\mathbf{E}_{g}X\|_{2}^{2}]=\langle\mathbf{K},\mathbf{E}_{g}^{T}\mathbf{E}_{g}\rangle=2\hskip-8.0pt\sum_{\{u,v\}\in\mathcal{P}}\hskip-10.0pt\mathcal{R}(u,v)+\hskip-20.0pt\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\hskip-20.0pt\sigma_{uv}\sigma_{uv^{\prime}}(\mathcal{R}(u,v)+\mathcal{R}(u,v^{\prime})-\mathcal{R}(v,v^{\prime})),$ (7)
@@
Several prior studies [31, 46, 59] have shown that certain structural patterns of adversarial edge perturbations are more impactful than others. The key distinction is that: previous approaches design perturbations primarily with respect to the graph topology, such as vertex degree, while our study advocates for jointly considering both the graph topology and the signal distribution. In earlier work, Waniek et al. [50] proposed the heuristic adversarial perturbation method DICE, which disconnects nodes within the same community and connect nodes across communities. Our theoretical analysis, along with the cSBM case study, provides a rigorous justification for this heuristic by explaining from a perspective of embedding perturbation why and when this strategy is efficient. We note that our framework is broadly applicable, as it neither relies on a specific graph filter nor requires ground-truth labels from downstream tasks, thereby providing insights into the structure of impactful perturbations across diverse graph learning settings.
@@ 5 Empirical Study on Adversarial Edge Perturbation
$\mathcal{P}^{*}=\operatorname*{arg\,max}_{|\mathcal{P}|=m}\mathbb{E}[\|\mathbf{E}_{g}X\|_{F}^{2}],\;\ \text{or }\mathcal{P}^{*}=\operatorname*{arg\,max}_{|\mathcal{P}|=m}\mathbb{E}[\|X^{(L)}-X^{(L)}_{p}\|_{F}^{2}].\vskip 0.0pt$
@@
$\hat{\mathcal{P}}=\operatorname*{arg\,max}_{|\mathcal{P}|=m}\langle\mathbf{K},\mathbf{E}_{g}^{T}\mathbf{E}_{g}\rangle.\vskip 0.0pt$ (8)
Here, the optimization objective involves the filter perturbation $\mathbf{E}_{g}$ and second-moment matrix $\mathbf{K}$. The filter perturbation $\mathbf{E}_{g}$ can be computed directly from its definition, $\mathbf{E}_{g}=g(\mathbf{S})-g(\mathbf{S}_{p})$, for any given filter and perturbation $\mathcal{P}$ (examples are provided in Appendix A.2). The second-moment matrix can be obtained using the empirical second-moment, i.e., $\mathbf{K}_{\text{samp}}=1/N\sum_{i=1}^{N}\mathbf{x}_{i}\mathbf{x}_{i}^{T}$, with alternative approaches discussed in Appendix C.3. The main obstacle to solving this optimization problem is the computational cost, as the complexity of brute-force searching is $O\left(n^{m}\right)$. To circumvent this, we consider relaxing discrete constraints in the problem and applying a projected gradient descent method developed by Xu et al. [54] for efficient adversarial attack in Prob-PGD. We present the algorithmic details in Appendix C.1. We also refer the reader to Appendix C.8, where we discuss how the proposed method can be adapted to large-scale datasets and present experimental results.
@@ 5.1 Adversarial embedding attack
Graph signals. Among the above datasets, only the ENZYMES dataset contains real-world measurement signals associated with their vertices, while the remaining five datasets only contain graph structures without signals on their vertices. For these graph-only datasets, we generate 100 synthetic signals from an arbitrary distribution, such as random Gaussian signals, signals from cSBM, or the smoothness distribution (see Appendix C.2 for a detailed description). The second-moment matrix $\mathbf{K}$ is obtained from computing the sample graph signals, i.e, using $\mathbf{K}_{\text{samp}}=1/100\sum_{i=1}^{100}\mathbf{x}_{i}\mathbf{x}_{i}^{T}$.
@@ 5.2 Adversarial attack on downstream tasks
Graph classification. We evaluate classification performance on two benchmark datasets: MUTAG (188 molecular graphs divided into two classes) and ENZYMES (600 protein graphs labeled by six enzyme classes). Both datasets provide numerical node features (e.g., atom types and physical measurements), from which we compute the second-moment matrix as $\mathbf{K}_{\text{samp}}=1/d\sum_{i=1}^{d}\mathbf{x}_{i}\mathbf{x}_{i}^{T}$. For each dataset, we use 80% of the graphs to pre-train 10 GCNNs and keep their parameters fixed during evaluation. To assess their prediction robustness, we perturb $5\%$ edges in test set graphs using different attack algorithms and measure classification accuracy before and after perturbation.
Node classification. We use the Cora citation network (2708 vertices belonging to seven categories) with sparse bag-of-words node features. We pre-train 10 GCNNs on the unperturbed graphs with 140 labeled nodes and keep the model parameters fixed during evaluation. To assess the robustness of these pre-trained models, we apply different algorithms to perturb 1000 edges to the graph and evaluate classification accuracy on test set nodes before and after perturbation. Additional experiments on heterophilic datasets are provided in Appendix C.7.
@@
[…]